OpenStudy (anonymous):

Have any of you heard of the Berry method of factoring? Does it always work?

OpenStudy (anonymous):

I might know it but not with that name

OpenStudy (anonymous):

Is it a variation of the grouping method or the trial and error method??

OpenStudy (anonymous):

Ok, say you have \[8s ^{2}+26s-45\] 1st, write (8s )(8s ) 2nd, multiply 8*(-45) which equals (-360) 3rd, take the factors of (-360) which add to +26 So, (+36)+(-10)=26 4th, (8s+36)(8s-10) 5th, factor=4(2s+9) 2(4s-5)

OpenStudy (anonymous):

if you are looking for the zeros of a quadratic equation, s= -9/2 s= 5/4

OpenStudy (anonymous):

I have seen this ... my teacher will not let us use it since each of the steps are not linked. For example, in your first step 8s times 8s is 64s^2 but you had 8s^2 when you started, you will divide out the extra 8 later, but we are not allowed to use Berry's method

OpenStudy (anonymous):

It is the fastest way that I have found to factor quadratics. However, when I had \[2x^2-7x-1=0\], using this method didn't work.

OpenStudy (anonymous):

I was wondering if that was related to the fact that the answer isn't a "real" number

OpenStudy (anonymous):

Yes, 2x^2 - 7x - 1 cannot be factored; therefore; Berry's factoring method would not work. 2x^2 - 7x - 1 is prime

OpenStudy (anonymous):

Oooh K. Thanks! Sometimes I forget to look and see if it even needs to be factored, I guess.